The following information applies to the questions displayed below.]

Trico Company set the following standard unit costs for its single product.

Direct materials (30 Ibs. @ $5.00 per Ib.) $ 150.00
Direct labor (7 hrs. @ $14 per hr.) 98.00
Factory overhead—variable (7 hrs. @ $7 per hr.) 49.00
Factory overhead—fixed (7 hrs. @ $9 per hr.) 63.00
Total standard cost $ 360.00


The predetermined overhead rate is based on a planned operating volume of 80% of the productive capacity of 61,000 units per quarter. The following flexible budget information is available.

Operating Levels
70% 80% 90%
Production in units 42,700 48,800 54,900
Standard direct labor hours 298,900 341,600 384,300
Budgeted overhead
Fixed factory overhead $ 3,074,400 $ 3,074,400 $ 3,074,400
Variable factory overhead $ 2,092,300 $ 2,391,200 $ 2,690,100


During the current quarter, the company operated at 90% of capacity and produced 54,900 units of product; actual direct labor totaled 316,300 hours. Units produced were assigned the following standard costs.

Direct materials (1,647,000 Ibs. @ $5.00 per Ib.) $ 8,235,000
Direct labor (384,300 hrs. @ $14 per hr.) 5,380,200
Factory overhead (384,300 hrs. @ $16 per hr.) 6,148,800
Total standard cost $ 19,764,000


Actual costs incurred during the current quarter follow.

Direct materials (1,364,000 Ibs. @ $7.80 per lb.) $ 10,639,200
Direct labor (316,300 hrs. @ $11.10 per hr.) 3,510,930
Fixed factory overhead costs 2,348,400
Variable factory overhead costs 2,742,200
Total actual costs $ 19,240,730

a) Compute the variable overhead spending and efficiency variances. (Round "cost per unit" and "rate per hour" answers to 2 decimal places.)

AH = Actual Hours
SH = Standard Hours
AVR = Actual Variable Rate
SVR = Standard Variable Rate

Actual Variable OH Cost -1 Flexible Budget -1 Standard Cost (VOH applied)
0
2
-1
$0
0

  
(b) Compute the fixed overhead spending and volume variances. (Round "cost per unit" and "rate per hour" answers to 2 decimal places.)

AH = Actual Hours
SH = Standard Hours
AFR = Actual Fixed Rate
SFR = Standard Fixed Rate

Actual Variable OH Cost Flexible Budget Standard Cost (VOH applied)
0
-1
$0
0

(c) Compute the total overhead controllable variance.
Answer #1

Solution a:

Variable Overhead Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH* AR = AH* SR = SH * SR =
316300 $8.67 $2,742,200.00 316300 $7.00 $2,214,100.00 384300 $7.00 $2,690,100.00
$528,100 U $476,000 F
Variable overhead rate Variance Variable overhead efficiency variance

Variable overhead spending variance = $528,100 U

Variable overhead efficiency variance = $476,000 F

Solution b:

Budgeted fixed overhead = 48800*7*$9 = $3,074,400

Actual fixed overhead = $2,348,400

Fixed overhead applied = Standard hours for actual production * SR = 54900*7*$9 = $3,458,700

Fixed overhead spending variance = Budgeted fixed overhead - Actual fixed overhead

= $3,074,400 - $2,348,400 = $726,000 F

Fixed overhead volume variance = Fixed overhead applied - Budgeted fixed overhead

= $3,458,700 - $3,074,400 = $384,300 F

Solution c:

Overhead controllable variance = Variable overhead spending variance + Variable overhead efficiency variance + Fixed overhead spending variance

= $528,100 U + $476,000 F + $726,000 F = $673,900 F